.60x+x=192 need help solving it
Found 2 solutions by HyperBrain, katya2:
Answer by HyperBrain(694)

Hi! Tnx 4 posting ur problem here.
Solution: (yahoo!)
I hope that I will not be needing to say too much words for you to understand.\\
60x+x=192
60x+1x=192
(60+1)x=192
61x=192
x=192/61=3.14754
Done!

Power up,
HyperBrain!

Answer by katya2(1)

60x+x
61x=192
192 : 61x
x=3.15